# TL-Training: A Task-Feature-Based Framework for Training Large Language Models in Tool Use

### v1

@@ Abstract
Large language models (LLMs) achieve remarkable advancements by leveraging tools to interact with external environments, a critical step toward generalized AI. However, the standard supervised fine-tuning (SFT) approach, which relies on large-scale datasets, often overlooks task-specific characteristics in tool use, leading to performance bottlenecks. To address this issue, we analyze three existing LLMs and uncover key insights: training data can inadvertently impede tool-use behavior, token importance is distributed unevenly, and errors in tool calls fall into a small set of distinct categories. Building on these findings, we propose TL-Training, a task-feature-based framework that mitigates the effects of suboptimal training data, dynamically adjusts token weights to prioritize key tokens during SFT, and incorporates a robust reward mechanism tailored to error categories, optimized through proximal policy optimization. We validate TL-Training by training CodeLLaMA-2-7B and evaluating it on four diverse open-source test sets. Our results demonstrate that the LLM trained by our method matches or surpasses both open- and closed-source LLMs in tool-use performance using only 1,217 training data points. Additionally, our method enhances robustness in noisy environments and improves general task performance, offering a scalable and efficient paradigm for tool-use training in LLMs. The code and data are available at https://github.com/Junjie-Ye/TL-Training.
@@ 1 Introduction
Large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023b; Bai et al., 2023) excel in natural language understanding and reasoning due to pre-training on extensive datasets (Chen et al., 2023; Ye et al., 2023). By incorporating tool-use capabilities, LLMs can extend beyond text generation to interact with the external environment, enabling tasks such as web searches and email management (Tang et al., 2023; Ye et al., 2024a). Furthermore, these capabilities are essential for addressing real-world user needs and advancing the development of general-purpose AI (Xi et al., 2023).
Current approaches to training LLMs for tool use rely heavily on large-scale datasets generated from trajectories of interactions with tools (Qin et al., 2024; Zhuang et al., 2023). Standard supervised fine-tuning (SFT) is then applied to pre-trained models (Touvron et al., 2023a, b). While effective in some cases, these methods overlook key task-specific characteristics, leading to performance bottlenecks. For instance, ToolLLaMA-2-7B-v2 (Qin et al., 2024) achieves only 80% of GPT-4’s performance on tool-use benchmarks (Ye et al., 2024a; Wu et al., 2024), indicating room for significant improvement.
To fill this gap, we conduct an in-depth analysis of three tool-using LLMs, uncovering several key phenomena. Notably, over 17% of the training data for RoTLLaMA contains tool-calling errors (Figure 1), primarily due to reliance on data from GPT series models, which are not entirely error-free with complex tools. Training on such flawed data can hinder model performance. Additionally, our tests of ToolLLaMA-2-7B-v2 and NexusRaven-13B-v2 (team, 2023) reveal that incorrect tool selections often share a common prefix with the correct ones (Table 1), and correcting initial erroneous tokens can lead to successful predictions. This suggests that certain tokens are more critical in tool selection. Moreover, the types of errors produced by tool calls are relatively limited (Figure 3), providing a foundation for targeted improvements across various error categories.
@@
Based on these insights, we propose TL-Training, a task-feature-based framework for training LLMs in tool use. TL-Training mitigates the negative impact of flawed training data by identifying erroneous interaction paths and excluding them from gradient updates. It prioritizes key tokens through adaptive weighting during SFT and incorporates tool feedback into a robust reward mechanism for reinforcement learning using the proximal policy optimization (PPO) (Schulman et al., 2017) algorithm.
We validate our approach by training CodeLLaMA-2-7B (Rozière et al., 2023) on a curated dataset of 1,217 tool-call trajectories generated with GPT-4o111https://platform.openai.com/docs/models/gpt-4o. Evaluations on four open-source test sets demonstrate that the model trained with TL-Training matches or surpasses the tool-use performance of leading open- and closed-source LLMs, despite requiring significantly less training data. Additionally, TL-Training improves robustness in noisy environments and enhances general task performance.
@@
We identify three key insights in tool-use training, including the impact of erroneous data, the uneven importance of tokens, and the constrained range of tool-calling error categories.
We propose TL-Training, a novel task-feature-based framework comprising of adverse effects mitigation, key tokens prioritization, and reinforcement learning to address misbehavior.
@@ 2 Preliminaries
For our analysis, we use the training set from RoTLLaMA, which includes 12,247 filtered multi-turn tool-call trajectories generated by GPT-4. Illustrated in Figure 1, 17% of these trajectories contain various errors in tool use, indicating that even advanced models like GPT-4 encounter challenges with complex tools. These erroneous trajectories pose a challenge for models trained through SFT, as they inherit these error patterns during learning. This predisposition to incorrect tool invocation highlights the need for more robust training methods to mitigate error propagation and improve overall model performance.
@@
We evaluate the performance of ToolLLaMA-2-7B-v2 and NexusRaven-13B-v2, which are built on LLaMA-2-7B (Touvron et al., 2023b) and CodeLLaMA-2-13B (Rozière et al., 2023), respectively, as representative tool-using LLMs. Our evaluation uses RoTBench (Clean) (Ye et al., 2024c), a manually labeled dataset for the single-turn tool-use task with standardized answers. The analysis focuses on errors related to tool selection, parameter identification, and content filling, with results shown in Table 1. We observe that when the model selects the wrong tool, it often chooses one with a prefix similar to the correct tool. By manually correcting the first incorrectly predicted token, the model can generate the correct one, suggesting that certain tokens are crucial for task success. Additionally, errors in parameter identification and content generation highlight areas where further fine-tuning is required to improve model performance.
Behavior Analysis
Our analysis of the model’s tool invocation behavior reveals that while successful invocations lack significant patterns, erroneous invocations fall into a finite set of distinguishable classes based on the tool, parameters, and contents. This provides a solid foundation for designing targeted enhancement strategies to address various error types.
@@ 3 Approaches
Building on the analysis in Section 2, we propose TL-Training, a novel training paradigm for LLMs in tool use. As shown in Figure 2, this paradigm incorporates three core techniques: mitigating the adverse effects of suboptimal data by preventing its back-propagation (Section 3.1), prioritizing key tokens using adaptive weight adjustments (Section 3.2), and implementing a reward mechanism tailored to tool invocation error categories to enable effective reinforcement learning (Section 3.3).222Theoretical proofs of the effectiveness of our proposed approaches is provided in Appendix A.
@@ 3.1 Mitigating Adverse Effects
Given a data sequence , we seek to identify the erroneous tool call trajectory . Directly determining whether a specific is correct is challenging. However, the feedback generated after each tool call contains structured error-reporting information, as summarized in Figure 3.333Specific examples can be found in Appendix C. We automate the identification of incorrect calls by sequentially analyzing to extract .
@@ 3.2 Prioritizing Key Tokens
Here, is the maximum adjustment multiplier, and is used to constrain the adjustment factor to lie within the range . The notation represents the size of the set.444Since always includes at least the first token of the tool name, we avoid any risk of dividing by zero.
@@ 3.3 Introducing a Reward Mechanism
This reward function addresses the different potential errors in LLM tool use, providing a structured scoring system to assess performance. Based on this reward function, we apply the PPO algorithm, which iteratively optimizes the model’s parameters to maximize these rewards as follows:
@@ 4.1 Dataset
To validate our approach, we construct a custom training set focused on multi-turn tool usage and evaluate it using four publicly available test sets. The corresponding statistics are provided in Table 2.
Training Data
While previous studies often construct over 100,000 data points for training (Qin et al., 2024), we deliberately limit our dataset size. Our main goal is to validate the effectiveness of our approach rather than to scale data volume. Surprisingly, the experimental results in Section 5 show that training on just 1,217 data points using our method matches or even exceeds the performance of leading LLMs.
Test Sets
To comprehensively evaluate LLM tool-use performance, we use four open-source tool usage test sets. ToolAlpaca (Tang et al., 2023), RoTBench (Ye et al., 2024c), and BFCL-v3 (Patil et al., 2023) are selected for single-turn tool use evaluations, while ToolEyes is used for multi-turn tool use assessment.555For consistency, we select the following versions for single-turn evaluations: ToolAlpaca (eval_real), RoTBench (Clean), and BFCL-v3 (executable).
@@
4.2 Baselines
Tool-Use LLMs
ToolLLaMA-2-7B and NexusRaven-2-13B are prominent tool-use LLMs, built on LLaMA-2-7B and CodeLLaMA-2-13B, respectively. These models are trained on a large volume of tool-use data, enhancing their ability to interact with tools. For example, ToolLLaMA-2-7B is trained on over 120,000 data points covering more than 16,000 tools using standard SFT, significantly boosting its tool-use capabilities.666Since NexusRaven-2-13B does not receive tool feedback during interactions, it is evaluated only on the single-turn tool-use datasets.
Open-Source LLMs
Among the existing open-source, general-purpose LLMs, ChatGLM-4-chat-9B (Zeng et al., 2024), Qwen-2-Instruct-7B (Yang et al., 2024a), LLaMA-3.1-Instruct-8B (Team, 2024a) and Qwen-2.5-Instruct-7B (Team, 2024b) have been specifically optimized for tool use, enabling them to interact with various tools to fulfill user needs.
Closed-Source LLMs
Our Model
@@ 4.3 Metrics
Given the distinctions between single-turn and multi-turn tool use, we design comprehensive metrics to evaluate the tool-use capabilities of LLMs.
Single-Turn Evaluation
For the evaluation of single-turn tool calls, where the original dataset provides a standard answer, we follow Ye et al. (2024c) and assess the model’s performance across three key areas:
Tool Selection (TS): Measures the model’s accuracy in selecting the tool specified by the standard answer.
Parameter Identification (PI): Evaluates the model’s ability to correctly select the tool and identify the relevant parameters required for invocation.
Content Filling (CF): Assesses the model’s capacity to complete the single-turn tool invocation, including selecting the correct tool, identifying relevant parameters, and filling in the appropriate values.
Multi-Turn Evaluation
For multi-turn tool use, where no standardized interaction path exists, we adapt the methods of Qin et al. (2024) and Ye et al. (2024a), and assess performance based on following metrics:
Documentation Understanding Error (DE): Represents the percentage of errors resulting from the model’s failure to interpret the tool documentation, encompassing tool hallucinations, parameter hallucinations, and missing necessary parameters.
Tool Call Error (CE): Denotes the proportion of errors arising from incorrect tool invocation, covering all error types except those classified as DE.
Valid Answers (VA): Evaluates the percentage of instances where the model delivers valid responses within nine turns, reflecting its ability to meet user needs effectively.
@@ 4.4 Implementation Details
Training
In the SFT stage, we use 1,217 constructed data samples, applying both the MAE and PKT strategies. We employ the AdamW optimizer (Loshchilov & Hutter, 2019) with cosine scheduling, setting the learning rate to 1e-6, a warmup rate of 0.01, and a batch size of 4, training for a total of 1 epoch. For the PKT strategy, is set to 9.
In the RL stage, we filter 1,194 entries from the constructed data and apply PPO with the reward function described in Section 3.3. The actor learning rate is set to 2e-6, the critic learning rate to 1e-6, and the batch size to 8, training for a total of 3 epochs.
Testing
For testing, we use the official prompt template for tool invocation,777Templates for each LLM are provided in Appendix B. and apply greedy search during inference to maximize each model’s tool-use capabilities.
@@ 5 Experiments
We compare the performance of nine LLMs, representing three different types, with our model across four tool-use evaluation sets (Section 5.1) and conduct detailed ablation experiments to validate the effectiveness of each component of our approach (Section 5.2).
@@ 5.1 Main Results
We evaluate the performance of various LLMs on three single-turn tool-use test sets and one multi-turn tool-use test set, with the results summarized in Table 3 and Table 4. Despite using the smallest model size and the least amount of training data, our approach achieves results comparable to the best-performing models. These findings demonstrate the potential for smaller, efficient models to excel in the tool use task, making advanced capabilities more accessible for resource-constrained environments.
@@
Results from three single-turn tool-use test sets demonstrate that TL-CodeLLaMA-2, with only 7B parameters and 1,217 training examples, surpasses all other open-source LLMs in overall task completion (i.e., CF). Remarkably, on the ToolAlpaca and BFCL-v3 datasets, TL-CodeLLaMA-2 outperforms GPT-4-turbo, the top-performing GPT family model, by an impressive 15.78% and 12.08%, respectively. Most notably, TL-CodeLLaMA-2 is the only model—among both open- and closed-source LLMs—that consistently exceeds the average performance across all three aspects of every dataset evaluated, highlighting the effectiveness of our approach in enhancing single-turn tool usage capabilities.

### v2

@@ Abstract
Large language models (LLMs) achieve remarkable advancements by leveraging tools to interact with environments, a critical step toward generalized AI. However, the standard supervised fine-tuning (SFT) approach, which relies on large-scale datasets, often overlooks task-specific characteristics in tool use, leading to performance bottlenecks. To address this issue, we analyze three existing LLMs and uncover key insights: training data can inadvertently impede tool-use behavior, token importance is distributed unevenly, and errors in tool calls fall into a small set of categories. Building on these findings, we propose TL-Training, a task-feature-based framework that mitigates the effects of suboptimal training data, dynamically adjusts token weights to prioritize key tokens during SFT, and incorporates a robust reward mechanism tailored to error categories, optimized through proximal policy optimization. We validate TL-Training by training CodeLLaMA-2-7B and evaluating it on four open-source test sets. Our results demonstrate that the LLM trained by our method matches or surpasses both open- and closed-source LLMs in tool-use performance using only 1,217 training data points. Additionally, our method enhances robustness in noisy environments and improves general task performance, offering a scalable and efficient paradigm for tool-use training in LLMs. Code and data are available at https://github.com/Junjie-Ye/TL-Training.
TL-Training: A Task-Feature-Based Framework for Training
Large Language Models in Tool Use
Junjie Ye1, Yilong Wu1, Sixian Li1, Yuming Yang1, Zhiheng Xi1, Tao Gui1,4, Qi Zhang1,4,5††thanks: Corresponding Author., Xuanjing Huang1,4,5, Peng Wang2, Zhongchao Shi2, Jianping Fan2, Zhengyin Du3 1Fudan University 2Lenovo Research, Beijing, China 3ByteDance Seed 4Shanghai Key Lab of Intelligent Information Processing 5Shanghai AI Laboratory jjye23@m.fudan.edu.cn, qz@fudan.edu.cn
@@ 1 Introduction
Large language models (LLMs) OpenAI (2023); Touvron et al. (2023); Bai et al. (2023) excel in natural language understanding due to pre-training on extensive datasets Chen et al. (2023); Ye et al. (2023). By incorporating tool-use capabilities, LLMs can extend beyond text generation to interact with the external environment, enabling tasks such as web searches and email management Tang et al. (2023); Ye et al. (2025). Furthermore, these capabilities are essential for addressing real-world user needs and advancing the development of general-purpose AI Xi et al. (2023).
Current approaches to training LLMs for tool use rely heavily on large-scale datasets generated from trajectories of interactions with tools Qin et al. (2024); Zhuang et al. (2023). Standard supervised fine-tuning (SFT) is then applied to pre-trained models. While effective in some cases, these methods overlook key task-specific characteristics, leading to performance bottlenecks. For instance, ToolLLaMA-2-7B-v2 Qin et al. (2024) achieves only 80% of GPT-4’s performance on tool-use benchmarks Ye et al. (2025); Wu et al. (2024), indicating room for significant improvement.
To fill this gap, we conduct an in-depth analysis of three tool-using LLMs, uncovering several key phenomena. Notably, over 17% of the training data for RoTLLaMA Ye et al. (2024b) contains tool-calling errors (Figure 1), primarily due to reliance on data from GPT series models, which are not entirely error-free with complex tools. Training on such flawed data can hinder model performance. Additionally, our tests of ToolLLaMA-2-7B-v2 and NexusRaven-13B-v2 team (2023) reveal that incorrect tool selections often share a common prefix with the correct ones (Table 1), and correcting initial erroneous tokens can lead to successful predictions. This suggests that certain tokens are more critical in tool selection. Moreover, the types of errors produced by tool calls are relatively limited (Figure 3), providing a foundation for targeted improvements across various error categories.
@@
Based on these insights, we propose TL-Training, a task-feature-based framework for training LLMs in tool use. TL-Training mitigates the negative impact of training data by identifying erroneous interaction paths and excluding them from gradient updates. It prioritizes key tokens through adaptive weighting during SFT and incorporates tool feedback into a robust reward mechanism for reinforcement learning using the proximal policy optimization (PPO) Schulman et al. (2017).
We validate our approach by training CodeLLaMA-2-7B Rozière et al. (2023) on a curated dataset of 1,217 tool-call trajectories generated with GPT-4o. Evaluations on four open-source test sets demonstrate that the model trained with TL-Training matches or surpasses the tool-use performance of leading open- and closed-source LLMs, despite requiring significantly less training data. Additionally, TL-Training improves robustness in noisy environments and enhances general task performance.
@@
We identify three key insights in tool use, including the impact of erroneous data, the uneven importance of tokens, and the constrained range of tool-calling error categories.
We propose TL-Training, a novel task-feature-based framework comprising of adverse effects mitigation, key tokens prioritization, and reinforcement learning for tool use.
@@
2 Related Works
Training LLMs for Tool Use
LLMs capable of utilizing external tools significantly enhance their ability to interact with dynamic environments and address user needs Qin et al. (2023). However, the diversity and complexity of real-world tools present significant challenges in training such models. Existing methods, such as SFT, rely on the generation of extensive datasets of tool interactions Song et al. (2023); Tang et al. (2023); Mekala et al. (2024), enabling models to learn tool functionalities, invoke appropriate tools, and process feedback. While effective, these methods are resource-intensive due to the large-scale data construction involved. To overcome these challenges, some studies have proposed encoding tool names as special tokens directly integrated into model training, embedding tool-specific knowledge into the model Hao et al. (2023). This approach has shown promise for existing tools but remains limited in its ability to adapt to newly introduced tools. Building on these findings, our work introduces a novel training paradigm for tool-use LLMs, addressing both efficiency and adaptability. By leveraging a compact dataset of 1,217 data points and incorporating three task-specific components, our approach achieves state-of-the-art performance while significantly reducing data requirements.
Evaluating LLMs in Tool Use
Evaluating LLMs’ tool-use capabilities is essential for understanding their effectiveness in diverse scenarios. A common evaluation method involves comparing predicted outputs with standard answers from a single turn of tool use Chen et al. (2024). However, in multi-turn interactions, the variability in invocation processes complicates the definition of a single standard path. To address this, evaluations increasingly consider multiple dimensions of tool-use processes and outcomes Ye et al. (2025). Beyond tool-use performance, researchers have also investigated robustness and safety in practical scenarios Ye et al. (2024b, a), which provide insights into how LLMs manage edge cases and avoid harmful outputs. In this paper, we evaluate LLMs across single-turn and multi-turn tool use to provide a more comprehensive assessment. Additionally, we analyze robustness to further demonstrate the superiority of our approach.
@@ 3 Preliminaries
For our analysis, we use the training set from RoTLLaMA, which includes 12,247 filtered multi-turn tool-call trajectories generated by GPT-4. Illustrated in Figure 1, 17% of these trajectories contain various errors, indicating that even advanced models like GPT-4 encounter challenges with complex tools. These erroneous trajectories pose a challenge for models trained through SFT, as they inherit these error patterns during learning. This predisposition to incorrect tool invocation highlights the need for more robust training methods to mitigate error propagation and improve overall model performance.
@@
We evaluate the performance of ToolLLaMA-2-7B-v2 and NexusRaven-13B-v2, which are built on LLaMA-2-7B Touvron et al. (2023) and CodeLLaMA-2-13B Rozière et al. (2023), respectively, as representative tool-using LLMs. Our evaluation uses RoTBench (Clean) Ye et al. (2024b), a manually labeled dataset for the single-turn tool-use task with standardized answers. The analysis focuses on errors related to tool selection, parameter identification, and content filling, with results shown in Table 1. We observe that when the model selects the wrong tool, it often chooses one with a prefix similar to the correct tool. By manually correcting the first incorrectly predicted token, the model can generate the correct one, suggesting that certain tokens are crucial for task success. Additionally, errors in parameter identification and content generation highlight areas where further training is required.
@@ 4 Approaches
Building on the analysis in Section 3, we propose TL-Training, a novel training paradigm for LLMs in tool use. As shown in Figure 2, this paradigm incorporates three core techniques: mitigating the adverse effects of suboptimal data by preventing its back-propagation (Section 4.1), prioritizing key tokens using adaptive weight adjustments (Section 4.2), and implementing a reward mechanism tailored to tool invocation error categories to enable effective reinforcement learning (Section 4.3).111Theoretical proofs of the effectiveness of our proposed approaches is provided in Appendix A.
@@ 4.1 Mitigating Adverse Effects
Given a data sequence , we seek to identify the erroneous tool call trajectory . Directly determining whether a specific is correct is challenging. However, the feedback generated after each tool call contains structured error-reporting information, as summarized in Figure 3.222Specific examples can be found in Appendix E. We automate the identification of incorrect calls by sequentially analyzing to extract .
@@ 4.2 Prioritizing Key Tokens
Here, is the maximum adjustment multiplier, and is used to constrain the adjustment factor to lie within the range . The notation represents the size of the set.333Since always includes at least the first token of the tool name, we avoid any risk of dividing by zero.
@@ 4.3 Introducing a Reward Mechanism
This reward function addresses the different potential errors in LLM tool use, providing a structured scoring system to assess performance. Based on this, we apply the PPO algorithm, which iteratively optimizes the model’s parameters to maximize these rewards as follows:
@@
5 Experimental Setup
5.1 Dataset
As shown in Table 2, to validate our approach, we construct a custom training set focused on multi-turn tool use and evaluate it using four publicly available test sets (i.e., ToolAlpaca Tang et al. (2023), RoTBench Ye et al. (2024b), BFCL-v3 Patil et al. (2024), and ToolEyes).444Details of the datasets can be found in Appendix B
@@ 5.1 Dataset
5.2 Baselines
We conduct a comprehensive comparison of ten LLMs from three different categories. These include ToolLLaMA-2-7B and NexusRaven-2-13B as tool-use LLMs; ChatGLM-4-chat-9B Zeng et al. (2024), Qwen-2-Instruct-7B Yang et al. (2024a), LLaMA-3.1-Instruct-8B Team (2024), and Qwen-2.5-Instruct-7B Yang et al. (2024b) as open-source LLMs; GPT-3.5-turbo, GPT-4o, and GPT-4-turbo as closed-source LLMs; and TL-CodeLLaMA-2, developed from CodeLLaMA-2-7B with the custom dataset.555Details of baselines can be found in Appendix C.
@@ 5.3 Metrics
For single-turn tool use, where the original dataset provides a standard answer, we follow Ye et al. (2024b) and assess the model’s performance across three key areas: 1) Tool Selection (TS) measures the model’s accuracy in selecting the tool specified by the standard answer; 2) Parameter Identification (PI) evaluates the model’s ability to correctly select the tool and identify the relevant parameters required for invocation; and 3) Content Filling (CF) assesses the model’s capacity to complete the single-turn tool invocation, including selecting the correct tool, identifying relevant parameters, and filling in the appropriate values.
For multi-turn tool use, where no standardized interaction path exists, we adapt the methods of Qin et al. (2024) and Ye et al. (2025), and assess performance based on following metrics: 1) Documentation Understanding Error (DE) represents the percentage of errors resulting from the model’s failure to interpret the tool documentation, encompassing tool hallucinations, parameter hallucinations, and missing necessary parameters; 2) Tool Call Error (CE) denotes the proportion of errors arising from incorrect tool invocation, covering all error types except those classified as DE; and 3) Valid Answers (VA) evaluates the percentage of instances where the model delivers valid responses within nine turns.
@@ 5.4 Implementation Details
In the SFT stage, we use 1,217 constructed data samples, applying both the MAE and PKT strategies. We employ the AdamW optimizer Loshchilov and Hutter (2019) with cosine scheduling, setting the learning rate to 1e-6, a warmup rate of 0.01, and a batch size of 4, training for a total of 1 epoch. For the PKT strategy, is set to 9. In the RL stage, we filter 1,194 entries from the constructed data and apply PPO with the reward function described in Section 4.3. The actor learning rate is set to 2e-6, the critic learning rate to 1e-6, and the batch size to 8, training for a total of 3 epochs. For testing, we use the official prompt template for tool invocation and apply greedy search.666Templates for each LLM are provided in Appendix D.
@@ 6.1 Main Results
We evaluate the performance of various LLMs on three single-turn tool-use test sets and one multi-turn tool-use test set, with the results summarized in Table 3 and Table 4.777Since NexusRaven-2-13B does not receive tool feedback during interactions, it is evaluated only on the single-turn tool-use datasets. Despite using the smallest model size and the least amount of training data, our approach achieves results comparable to the best-performing models. These findings demonstrate the potential for smaller, efficient models to excel in the tool use task, making advanced capabilities more accessible for resource-constrained environments.
@@
Results from three single-turn tool-use test sets demonstrate that TL-CodeLLaMA-2, with only 7B parameters and 1,217 training examples, surpasses all other open-source LLMs in overall task completion (i.e., CF). Remarkably, on the ToolAlpaca and BFCL-v3 datasets, TL-CodeLLaMA-2 outperforms GPT-4-turbo, the top-performing GPT family model, by an impressive 15.78% and 12.08%, respectively. Most notably, TL-CodeLLaMA-2 is the only model that consistently exceeds the average performance across all three aspects of every dataset evaluated, highlighting the effectiveness of our approach in enhancing single-turn tool usage capabilities.